OpenStudy (anonymous):

Not sure what i'm doing wrong here. I had to calculate the Cs of 5 different metals. As of right now I have only done zinc, but I got 13.22J/gCas my Cs of it and the literature value is WAY off. It's .39 J/gC at 25 degree C.

OpenStudy (anonymous):

We had 50 mL of water. a 15.825g sample of zinc. Ti = 20.40 degree C and the Tf = 19.70 degree C. so this is what I did \[-(50g \times 4.184 J/gC \times -.70) = (15.825g \times C _{s} of metal \times .70)\]